OpenStudy (anonymous):

use the echelon method to solve the following system of three equations 5x-2y+5z=59 3x-2y-3z=-1 x-y+2z=12

OpenStudy (anonymous):

Hey sorry for not responding sooner (it takes me a while to do these). Anyway I would first set up the matrix like so: [5 -2 5|59] [3 -2 -3|-1] [1 -1 2|12] If you swap the first with the third row you have this: [1 -1 2|12] [3 -2 -3|-1] [5 -2 5|59] Next let your new Row2=row2-3*(row1) and let your new Row 3 = row3-5*(row1). Then you will have: [1 -1 2|12] [0 1 -9|-37] [0 3 -5|-1] Next make your new Row3=row3-3*(row2). Then you should have: [1 -1 2|12] [0 1 -9|-37] [0 0 22|110] Then for your new Row3 multiply 1/22 with your old row3. So you have: [1 -1 2|12] [0 1 -9|-37] [0 0 1| 5] Now you have enough information to solve the system of equations. z=5 y-9z=-37 x-y+2z=12 I leave it to you to finish. :D Did that help?